Grade 12General Physics

Two identical metallic spheres having unequal opposite charges are placed at a distance of 0.50m apart in air. After bringing them in contact with each other they are again placed at the same distance apart. Now the force of repulsion between them is 0.108N. Calculate the final charge on each of them?

Ranbir
9 Years agoGrade 12

1 Answer

kalavena srikanth
9 Years ago
F=kq1q2/d^2. Where d is distance between two bodies. K=9×10^9,Nm^2/coulomb^2. By substitute we get F=36×10^9q^2N. q1=q2=q since two bodies are identical. given F=.108N by substitute. We get q^2=3×10^_12 then q=+1.732×10^-6and q=-1.732×10^-6